# Annabelle Michael Carrell    Albert Gong    Abhishek Shetty    Raaz Dwivedi    Lester Mackey

### v5

@@ B.1.2 Sub-Gaussianity of subsampling
Fix any vector $\boldsymbol{u}\in\mathbb{R}^{n}$, and let $J_{1},\dots,J_{n_{\textup{out}}}$ be the random indices in $[n]$ selected by uniform subsampling. Since $\boldsymbol{u}^{\top}\mathbf{K}(\boldsymbol{p}_{\textup{in}}-\boldsymbol{p}_{\textup{out}})=\frac{1}{n_{\textup{out}}}\sum_{i=1}^{n_{\textup{out}}}\boldsymbol{u}^{\top}\mathbf{K}(\boldsymbol{p}_{\textup{in}}-\boldsymbol{e}_{J_{i}})$ is an average of mean-centered scalars drawn without replacement and satisfying
@@ B.2 KH$(\delta)$
Suppose $n_{\textup{in}}\geq 2$. In the notation of Alg. B.1, on a common event $\mathcal{E}$ of probability at least $21-\delta/$ 2, for all $i\in[n_{\textup{in}}/2]$, $\frac{1}{2i}\widetilde{\psi}_{i}$ is $(\mathbf{k},\nu_{i})$-sub-Gaussian with
@@
Prop. B.2 and the triangle inequality imply that $(\mathbb{P}_{\textup{in}}-\mathbb{P}_{\textup{out}})\mathbf{k}=\frac{1}{n_{\textup{in}}}\psi_{n_{\textup{in}}/2}$ is $(\mathbf{k},\nu)$-sub-Gaussian on $\mathcal{E}$ with
@@
For each $i$, the SBHW sub-Gaussian constants 70 satisfy
@@
Suppose $\sigma_{i-1}^{2}\leq c_{i-1}$. Then $\sigma_{i}^{2}=g(\sigma_{i-1}^{2})$ for $g(x)=x+\mathfrak{b}_{i}^{2}(1-x/r_{i})_{+}$. Note that the slope of $g$ is $1-\mathfrak{b}_{i}^{2}/r_{i}$ for $x<r_{i}$ and $1$ for $x>r_{i}$. If $1-\mathfrak{b}_{i}^{2}/r_{i}\geq 0$, then $g$ is increasing and its maximum value over $[0,c_{i}]$ is at $c_{i}$. If, on the other hand, $01-\mathfrak{b}_{i}^{2}/r_{i}<$ 0, then $g$ first decreases and then increases so its maximum value over $[0,c_{i}]$ is either at 00 or at $c$. Since $c_{i}\geq\max(r_{i},c_{i-1})$, $\sigma_{i}^{2}\leq\max(g(0),g(c_{i}))=\max(\mathfrak{b}_{i}^{2},c_{i})=c_{i}$. The proof is complete. ∎ Invoking Lem. B.1, the assumption $n_{\textup{in}}\geq 2$, and the fact that $\delta\mapsto\frac{\frac{1}{2}+\log(2/\delta)}{\log(2/\delta)}$ is increasing on $(0,1]$, we find that
@@
The first inequality in 72 and the definition 70 further imply that
@@
Hence, by Dwivedi & Mackey (2024, Thm. 3(iii)), for each $i\in[n_{\textup{in}}/2]$, the vector $\widetilde{\psi}_{i}$ of KH$(\delta)$ coincides with the vector $\psi_{i}$ of SBHW on a common event $\mathcal{E}$ of probability at least $21-\delta/$ 2. Therefore, each $\frac{1}{2i}\widetilde{\psi}_{i}$ is $(\mathbf{k},\frac{1}{2i}\sigma_{i})$-sub-Gaussian on $\mathcal{E}$, implying the result.
@@ B.3 LKH$(\delta)$
       $(\mathfrak{a}_{i},\sigma_{i})\leftarrow$ get_swap_params($\sigma_{i-1},\mathfrak{b}_{i},\delta_{i}$)
@@
function get_swap_params($\sigma,\mathfrak{b},\delta$):
@@
Suppose $n_{\textup{in}}\geq 2$. In the notation of Alg. B.2, on a common event $\mathcal{E}$ of probability at least $21-\delta/$ 2, for all $i\in[n_{\textup{in}}/2]$, $\frac{1}{2i}\widetilde{\psi}_{i}$ is $(\mathbf{k},\nu_{i})$-sub-Gaussian with $\mathbf{k}(\boldsymbol{x},\boldsymbol{y})=\langle{\boldsymbol{x}},{\boldsymbol{y}}\rangle$ and
@@
Dwivedi & Mackey (2024, Thm. 3(iii)) implies that, for each $i\in[n_{\textup{in}}/2]$, the vector $\widetilde{\psi}_{i}$ of LKH$(\delta)$ coincides with the vector $\psi_{i}$ of SBHW on a common event $\mathcal{E}$ of probability at least $21-\delta/$ 2. Therefore, each $\frac{1}{2i}\widetilde{\psi}_{i}$ is $(\mathbf{k},\frac{1}{2i}\sigma_{i})$-sub-Gaussian on $\mathcal{E}$. Finally, Dwivedi & Mackey (2024, (46)) shows that $\sigma_{i}\leq\nu_{i}$ for each $i\in[n_{\textup{in}}/2]$, yielding the result.
@@ B.4 RKH$(\delta)$
on an event $\mathcal{E}$ of probability at least $21-\delta/$ 2.
@@
Let $c=2\min(\max_{\boldsymbol{x}\in\mathcal{X}_{\textup{in}}}\sqrt{\mathbf{k}(\boldsymbol{x},\boldsymbol{x})},\max_{\boldsymbol{x}\in\mathcal{X}_{\textup{in}}}\operatorname{MMD}_{\mathbf{k}}(\boldsymbol{\delta}_{\boldsymbol{x}},\mathbb{P}_{\textup{in}}))$, and, for each $\ell\in[m]$, let $\widetilde{\psi}^{(\ell)}$ represent the vector $\widetilde{\psi}_{n_{\textup{in}}/2^{\ell}}$ produced at the end of the $\ell$-th call to KH$(\delta)$. By the proof of Prop. B.2 and the union bound, on an event $\mathcal{E}$ of probability at least $21-\delta/$ 2, $(\widetilde{\psi}^{(\ell)})_{\ell\in[m]}=(\psi^{(\ell)})_{\ell\in[m]}$, where each $\frac{2^{\ell-1}}{n_{\textup{in}}}\psi^{(\ell)}$ is $(\mathbf{k},\nu^{(\ell)})$-sub-Gaussian given $(\psi^{(j)})_{j\in[\ell-1]}$ for
@@ B.5 KH-Compress$(\delta)$
            $\widetilde{\mathcal{S}_{i}}\leftarrow$ compress($\mathcal{S}_{i}$)                  // return coresets of size $2^{\mathfrak{g}}\cdot\sqrt{\frac{\mathopen{}\mathclose{{}\left|\mathcal{S}}\right|}{4}}$
@@
return compress($\mathcal{X}_{\textup{in}}$)                  // coreset of size $n_{\textup{out}}=2^{\mathfrak{g}}\sqrt{n_{\textup{in}}}$
@@
on an event $\mathcal{E}$ of probability at least $21-\delta/$ 2.
@@
Now fix any SPSD $\mathbf{K}$ and any kernel $\mathbf{k}$ that generates $\mathbf{K}$. By Lem. A.1, we have that $\boldsymbol{p}_{\textup{in}}-\boldsymbol{p}_{\textup{out}}$ is $(\mathbf{K},\nu)$-sub-Gaussian on $\mathcal{E}$ and hence that KH-Compress$(\delta)\in\mathcal{G}_{\nu,\delta}(\mathbf{K})$. In addition, $\nu=O(\frac{\sqrt{\log(n_{\textup{out}})\log(n_{\textup{out}}/\delta)}}{n_{\textup{out}}})$ when $n_{\textup{out}}\geq\sqrt{n_{\textup{in}}}$. Furthermore, Shetty et al. (2022, Rem. 1) implies that KH-Compress$(\delta)$ has a runtime less than $4^{\mathfrak{g}+1}n_{\textup{in}}\mathopen{}\mathclose{{}\left(\log_{4}(n_{\textup{in}})-\mathfrak{g}}\right)=4n_{\textup{out}}^{2}\log_{2}(n_{\textup{in}}/n_{\textup{out}})=O(n_{\textup{out}}^{2})$ when $n_{\textup{out}}\geq\sqrt{n_{\textup{in}}}$.
@@ B.6 GS-Thin
$\boldsymbol{z}\leftarrow$ kernel_gs_walk($\mathcal{X}_{\textup{in}}$)
@@
             // Update set of active coordinates by removing smallest index set to ±$\pm 1$
@@
             $\boldsymbol{u}_{t}\leftarrow\mathop{\mathrm{missing}}{argmin}_{\boldsymbol{u}\in\mathbb{R}^{n_{\textup{in}}/2}}\boldsymbol{u}^{\top}\mathbf{Q}\boldsymbol{u}$ subject to $\boldsymbol{u}_{p^{\prime}}=1$ and $\boldsymbol{u}_{i}=0$ for all $i\notin\mathcal{A}^{\prime}$,
@@
$\boldsymbol{z}\leftarrow$ kernel_gs_walk_cubic($\mathcal{X}_{\textup{in}}$)
@@
function kernel_gs_walk_cubic($\mathopen{}\mathclose{{}\left(x_{i}}\right)_{i=1}^{n_{\textup{in}}}$):
@@
             // Update set of active coordinates by removing smallest index set to ±$\pm 1$
@@
Let $\boldsymbol{z}_{1},\boldsymbol{z}_{2},\dots$ be the fractional assignment sequence generated by kernel_gs_walk($(\boldsymbol{x}_{i})_{i=1}^{n_{\textup{in}}}$) in Alg. B.6 and $\boldsymbol{z}_{1}^{\prime},\boldsymbol{z}_{2}^{\prime},\dots$ be the fractional assignment sequence generated by kernel_gs_walk_cubic($(\boldsymbol{x}_{i})_{i=1}^{n_{\textup{in}}}$) in Alg. B.7 with an identical source of randomness. If the pairwise difference matrix
@@
Simplifying the above using the fact that $n_{\textup{out}}=n_{\textup{in}}/2^{m}$ yields our desired result 83.
@@
$\boldsymbol{u}_{t}\leftarrow\mathop{\mathrm{missing}}{argmin}_{\boldsymbol{u}\in\mathbb{R}^{n_{\textup{in}}/2}}\boldsymbol{u}^{\top}\mathbf{Q}\boldsymbol{u}$ (91)
@@
$\boldsymbol{u}_{t}\leftarrow\mathop{\mathrm{missing}}{argmin}_{\boldsymbol{u}\in\mathbb{R}^{n}}\boldsymbol{u}^{\top}\mathbf{Q}\boldsymbol{u}\quad\text{subject to}\quad\boldsymbol{u}_{p^{\prime}}=1\quad\text{and}\quad\boldsymbol{u}_{i}=0\quad\text{for all}\quad i\notin\mathcal{A}^{\prime}.$ (94)
@@
Hence, if we already have access to a matrix $\mathbf{C}=(\mathbf{Q}_{\mathcal{A}_{1}\times\mathcal{A}_{1}})^{-1}$, we can compute $\mathbf{D}$ by dropping the row and column of $\mathbf{C}$ corresponding to $i$ and then compute $(\mathbf{Q}_{\mathcal{A}_{2}\times\mathcal{A}_{2}})^{-1}$ using 99. Since in Alg. B.7 we begin by explicitly computing the inverse of $\mathbf{Q}_{\mathcal{A}^{\prime}\times\mathcal{A}^{\prime}}$, the update step in Alg. B.7 maintains the required inverse and thus its partial assignment updates match those of Alg. B.6.
@@
$\sum_{i=0}^{m-1}r_{\mathrm{H}}(n_{\textup{in}}/2^{i})=\sum_{i=0}^{m-1}C(n_{\textup{in}}/2^{i})^{3}=O(n_{\textup{in}}^{3}).$ (101)
@@ B.7 GS-Compress
            $\widetilde{\mathcal{S}_{i}}\leftarrow$ compress($\mathcal{S}_{i}$)                  // return coresets of size $2^{\mathfrak{g}}\cdot\sqrt{\frac{\mathopen{}\mathclose{{}\left|\mathcal{S}}\right|}{4}}$
@@
return compress($\mathcal{X}_{\textup{in}}$)                  // coreset of size $n_{\textup{out}}=2^{\mathfrak{g}}\sqrt{n_{\textup{in}}}$
@@ C.1 Proof of MMD bound 12
Without loss of generality, we suppose that $r\leq\mathrm{rank}({\mathbf{K}})$. Let $\mathbf{V}\boldsymbol{\Lambda}\mathbf{V}^{\top}$ be an eigendecomposition of $\mathbf{K}$ with orthonormal $\mathbf{V}\in\mathbb{R}^{n\times n}$ and diagonal $\boldsymbol{\Lambda}=\mathop{\mathrm{missing}}{diag}(\lambda_{1},\cdots,\lambda_{n})\in{}^{n\times n}$. Let $\mathbf{V}_{r}$ represent the first $r$ columns of $\mathbf{V}$, and let $\mathbf{V}_{-r}$ represent the last $n-r$ columns of $\mathbf{V}$. Introduce the shorthand
$\boldsymbol{w}\triangleq\boldsymbol{p}_{\textup{in}}-\boldsymbol{p}_{\textup{out}}\in{}^{n}\quad\text{and}\quad\boldsymbol{\Phi}\triangleq\mathbf{V}\boldsymbol{\Lambda}^{1/2}\mathbf{V}^{\top}\in{}^{n\times n}.$ (106)
@@
which when combined with 109 yield the advertised claim 12 on the squared MMD.
C.1.1 Proof of 110: Bounding $\|{\mathbf{V}_{r}^{\top}\boldsymbol{\Phi}^{\top}\boldsymbol{w}}\|_{2}^{2}$
@@
Rearranging terms yields the claimed bound 112. ∎
@@
Now fix any $\boldsymbol{u}\in\mathcal{C}_{\varepsilon,r}$ and let $\boldsymbol{\Lambda}_{r}=\mathop{\mathrm{missing}}{diag}(\lambda_{1},\dots,\lambda_{r})$. Using 106 and 107, we have
@@
$(1+\frac{2}{\varepsilon})^{r}\exp(\frac{-(1-\varepsilon)^{2}\alpha^{2}}{2\nu^{2}})\leq\delta^{\prime}\quad\text{if}\quad\alpha\geq\frac{\nu\sqrt{2}}{1-\varepsilon}\sqrt{\log(\frac{1}{\delta^{\prime}})+r\log(1+\frac{2}{\varepsilon})}$ (131)
@@
C.1.2 Proof of 111: Bounding $\|{\mathbf{V}_{-r}^{\top}\boldsymbol{\Phi}^{\top}\boldsymbol{w}}\|_{2}^{2}$
@@
we have, for $\boldsymbol{\Lambda}_{-r}\triangleq\mathop{\mathrm{missing}}{diag}(\lambda_{r+1},\cdots,\lambda_{n})$ and $\lambda_{\max}$ the maximum eigenvalue of a SPSD matrix,

### v7

@@ B.1.2 Sub-Gaussianity of subsampling
[…]
@@ B.2 KH$(\delta)$
Suppose $n_{\textup{in}}\geq 2$. In the notation of Alg. B.1, on a common event $\mathcal{E}$ of probability at least $1-\delta/2$, for all $i\in[n_{\textup{in}}/2]$, $\frac{1}{2i}\widetilde{\psi}_{i}$ is $(\mathbf{k},\nu_{i})$-sub-Gaussian with
@@
Prop. B.2 and the triangle inequality imply that $(\mathbb{P}_{\textup{in}}-\mathbb{P}_{\textup{out}})\mathbf{k}=\frac{1}{n_{\textup{in}}}\widetilde{\psi}_{n_{\textup{in}}/2}$ is $(\mathbf{k},\nu)$-sub-Gaussian on $\mathcal{E}$ with
@@
For each $i$, the SBHW sub-Gaussian constants 67 satisfy
@@
Suppose $\sigma_{i-1}^{2}\leq c_{i-1}$. Then $\sigma_{i}^{2}=g(\sigma_{i-1}^{2})$ for $g(x)=x+\mathfrak{b}_{i}^{2}(1-x/r_{i})_{+}$. Note that the slope of $g$ is $1-\mathfrak{b}_{i}^{2}/r_{i}$ for $x<r_{i}$ and 111 for $x>r_{i}$. If $1-\mathfrak{b}_{i}^{2}/r_{i}\geq 0$, then $g$ is increasing and its maximum value over $[0,c_{i}]$ is at $c_{i}$. If, on the other hand, $1-\mathfrak{b}_{i}^{2}/r_{i}<0$, then $g$ first decreases and then increases so its maximum value over $[0,c_{i}]$ is either at 0 or at $c_{i}$. Since $c_{i}\geq\max(r_{i},c_{i-1})$, $\sigma_{i}^{2}\leq\max(g(0),g(c_{i}))=\max(\mathfrak{b}_{i}^{2},c_{i})=c_{i}$. The proof is complete. ∎ Invoking Lem. B.1, the assumption $n_{\textup{in}}\geq 2$, and the fact that $\delta\mapsto\frac{\frac{1}{2}+\log(2/\delta)}{\log(2/\delta)}$ is increasing on (0,1](0,1]( 0 , 1 ], we find that
@@
The first inequality in 69 and the definition 67 further imply that
@@
Hence, by Dwivedi & Mackey (2024, Thm. 3(iii)), for each $i\in[n_{\textup{in}}/2]$, the vector $\widetilde{\psi}_{i}$ of KH$(\delta)$ coincides with the vector $\psi_{i}$ of SBHW on a common event $\mathcal{E}$ of probability at least $1-\delta/2$. Therefore, each $\frac{1}{2i}\widetilde{\psi}_{i}$ is $(\mathbf{k},\frac{1}{2i}\sigma_{i})$-sub-Gaussian on $\mathcal{E}$, implying the result.
@@ B.3 LKH$(\delta)$
[…]
@@
[…]
@@
Suppose $n_{\textup{in}}\geq 2$. In the notation of Alg. B.2, on a common event $\mathcal{E}$ of probability at least $1-\delta/2$, for all $i\in[n_{\textup{in}}/2]$, $\frac{1}{2i}\widetilde{\psi}_{i}$ is $(\mathbf{k},\nu_{i})$-sub-Gaussian with $\mathbf{k}(\boldsymbol{x},\boldsymbol{y})=\langle{\boldsymbol{x}},{\boldsymbol{y}}\rangle$ and
@@
Dwivedi & Mackey (2024, Thm. 3(iii)) implies that, for each $i\in[n_{\textup{in}}/2]$, the vector $\widetilde{\psi}_{i}$ of LKH$(\delta)$ coincides with the vector $\psi_{i}$ of SBHW on a common event $\mathcal{E}$ of probability at least $1-\delta/2$. Therefore, each $\frac{1}{2i}\widetilde{\psi}_{i}$ is $(\mathbf{k},\frac{1}{2i}\sigma_{i})$-sub-Gaussian on $\mathcal{E}$. Finally, Dwivedi & Mackey (2024, (46)) shows that $\sigma_{i}\leq\nu_{i}$ for each $i\in[n_{\textup{in}}/2]$, yielding the result.
@@ B.4 RKH$(\delta)$
on an event $\mathcal{E}$ of probability at least $1-\delta/2$.
@@
Let $c=2\min(\max_{\boldsymbol{x}\in\mathcal{X}_{\textup{in}}}\sqrt{\mathbf{k}(\boldsymbol{x},\boldsymbol{x})},\max_{\boldsymbol{x}\in\mathcal{X}_{\textup{in}}}\operatorname{MMD}_{\mathbf{k}}(\boldsymbol{\delta}_{\boldsymbol{x}},\mathbb{P}_{\textup{in}}))$, and, for each $\ell\in[m]$, let $\widetilde{\psi}^{(\ell)}$ represent the vector $\widetilde{\psi}_{n_{\textup{in}}/2^{\ell}}$ produced at the end of the $\ell$-th call to KH$(\delta)$. By the proof of Prop. B.2 and the union bound, on an event $\mathcal{E}$ of probability at least $1-\delta/2$, $(\widetilde{\psi}^{(\ell)})_{\ell\in[m]}=(\psi^{(\ell)})_{\ell\in[m]}$, where each $\frac{2^{\ell-1}}{n_{\textup{in}}}\psi^{(\ell)}$ is $(\mathbf{k},\nu^{(\ell)})$-sub-Gaussian given $(\psi^{(j)})_{j\in[\ell-1]}$ for
@@ B.5 KH-Compress$(\delta)$
[…]
@@
[…]
@@
on an event $\mathcal{E}$ of probability at least $1-\delta/2$.
@@
Prop. B.5 and Lem. A.1 imply that KH-Compress$(\delta)\in\mathcal{G}_{\nu,\delta}(\mathbf{K})$ for any kernel $\mathbf{k}$ that generates $\mathbf{K}$. In addition, $\nu=O(\frac{\sqrt{\log(n_{\textup{out}})\log(n_{\textup{out}}/\delta)}}{n_{\textup{out}}})$ when $n_{\textup{out}}\geq\sqrt{n_{\textup{in}}}$. Furthermore, Shetty et al. (2022, Rem. 1) implies that KH-Compress$(\delta)$ has a runtime less than $4^{\mathfrak{g}+1}n_{\textup{in}}\mathopen{}\mathclose{{\left(\log_{4}(n_{\textup{in}})-\mathfrak{g}}}\right)=4n_{\textup{out}}^{2}\log_{2}(n_{\textup{in}}/n_{\textup{out}})=O(n_{\textup{out}}^{2})$ when $n_{\textup{out}}\geq\sqrt{n_{\textup{in}}}$.
@@ B.6 GS-Thin
[…]
@@
     // Update set of active coordinates by removing smallest index set to ±1$\pm$ 1± 1
@@
     $\boldsymbol{u}_{t}\leftarrow\mathop{\mathrm{argmin}}_{\boldsymbol{u}\in\mathbb{R}^{n_{\textup{in}}/2}}\boldsymbol{u}^{\top}\mathbf{Q}\boldsymbol{u}$ subject to $\boldsymbol{u}_{p^{\prime}}=1$ and $\boldsymbol{u}_{i}=0$ for all $i\notin\mathcal{A}^{\prime}$,
@@
[…]
@@
[…]
@@
     // Update set of active coordinates by removing smallest index set to ±1$\pm$ 1± 1
@@
[…]
@@
Simplifying the above using the fact that $n_{\textup{out}}=n_{\textup{in}}/2^{m}$ yields our desired result 80.
@@
$\boldsymbol{u}_{t}\leftarrow\mathop{\mathrm{argmin}}_{\boldsymbol{u}\in\mathbb{R}^{n_{\textup{in}}/2}}\boldsymbol{u}^{\top}\mathbf{Q}\boldsymbol{u}$ (88)
@@
$\boldsymbol{u}_{t}\leftarrow\mathop{\mathrm{argmin}}_{\boldsymbol{u}\in\mathbb{R}^{n}}\boldsymbol{u}^{\top}\mathbf{Q}\boldsymbol{u}\quad\text{subject to}\quad\boldsymbol{u}_{p^{\prime}}=1\quad\text{and}\quad\boldsymbol{u}_{i}=0\quad\text{for all}\quad i\notin\mathcal{A}^{\prime}.$ (91)
@@
Hence, if we already have access to a matrix $\mathbf{C}=(\mathbf{Q}_{\mathcal{A}_{1}\times\mathcal{A}_{1}})^{-1}$, we can compute $\mathbf{D}$ by dropping the row and column of $\mathbf{C}$ corresponding to $i$ and then compute $(\mathbf{Q}_{\mathcal{A}_{2}\times\mathcal{A}_{2}})^{-1}$ using 96. Since in Alg. B.7 we begin by explicitly computing the inverse of $\mathbf{Q}_{\mathcal{A}^{\prime}\times\mathcal{A}^{\prime}}$, the update step in Alg. B.7 maintains the required inverse and thus its partial assignment updates match those of Alg. B.6.
@@
$\sum_{i=0}^{m-1}r_{\mathrm{H}}(n_{\textup{in}}/2^{i})\leq\sum_{i=0}^{m-1}C(n_{\textup{in}}/2^{i})^{3}=O(n_{\textup{in}}^{3}).$ (98)
@@ B.7 GS-Compress
[…]
@@
  return GS-Halve-Cubic$(\widetilde{\mathcal{S}},\mathbf{k})$            // coreset of size $2^{\mathfrak{g}}\sqrt{\mathopen{}\mathclose{{\left|\mathcal{S}}}\right|}$
@@ C.1 Proof of MMD bound 12
Without loss of generality, we suppose that $r\leq\mathrm{rank}({\mathbf{K}})$. Let $\mathbf{V}\boldsymbol{\Lambda}\mathbf{V}^{\top}$ be an eigendecomposition of $\mathbf{K}$ with orthonormal $\mathbf{V}\in\mathbb{R}^{n\times n}$ and diagonal $\boldsymbol{\Lambda}=\mathop{\mathrm{diag}}(\lambda_{1},\cdots,\lambda_{n})\in\real^{n\times n}$. Let $\mathbf{V}_{r}$ represent the first $r$ columns of $\mathbf{V}$, and let $\mathbf{V}_{-r}$ represent the last $n-r$ columns of $\mathbf{V}$. Introduce the shorthand
$\boldsymbol{w}\triangleq\boldsymbol{p}_{\textup{in}}-\boldsymbol{p}_{\textup{out}}\in\real^{n}\quad\text{and}\quad\boldsymbol{\Phi}\triangleq\mathbf{V}\boldsymbol{\Lambda}^{1/2}\mathbf{V}^{\top}\in\real^{n\times n}.$ (103)
@@
which when combined with 106 yield the advertised claim 12 on the squared MMD.
C.1.1 Proof of 107: Bounding $\|{\mathbf{V}_{r}^{\top}\boldsymbol{\Phi}^{\top}\boldsymbol{w}}\|_{2}^{2}$
@@
Rearranging terms yields the claimed bound 109. ∎
@@
Now fix any $\boldsymbol{u}\in\mathcal{C}_{\varepsilon,r}$ and let $\boldsymbol{\Lambda}_{r}=\mathop{\mathrm{diag}}(\lambda_{1},\dots,\lambda_{r})$. Using 103 and 104, we have
@@
$(1+\frac{2}{\varepsilon})^{r}\exp(\frac{-(1-\varepsilon)^{2}\alpha^{2}}{2\nu^{2}})\leq\delta^{\prime}\quad\text{if}\quad\alpha\geq\frac{\nu\sqrt{2}}{1-\varepsilon}\sqrt{\log(\frac{1}{\delta^{\prime}})+r\log(1+\frac{2}{\varepsilon})}.$ (128)
@@
C.1.2 Proof of 108: Bounding $\|{\mathbf{V}_{-r}^{\top}\boldsymbol{\Phi}^{\top}\boldsymbol{w}}\|_{2}^{2}$
@@
we have, for $\boldsymbol{\Lambda}_{-r}\triangleq\mathop{\mathrm{diag}}(\lambda_{r+1},\cdots,\lambda_{n})$ and $\lambda_{\max}$ the maximum eigenvalue of a SPSD matrix,